# Hierarchical Insights: Exploiting Structural Similarities for Reliable 3D Semantic Segmentation

### v2

@@ III-A Data and Label Hierarchy
The label hierarchy serves to group semantically similar classes together. Our proposed label hierarchy for the SemanticKITTI dataset is illustrated in Fig. 2. Each original label $y$ is associated with multiple superclasses $\mathcal{S}$ on a level $l$. The dataset’s original ground truth labels represent the leaf nodes ($l=0$). Superclasses are structured as follows: the second level aligns with SemanticKITTI’s suggestion ($l=1$), and the third level introduces binary classes, namely static and dynamic ($l=2$). The root node encompasses only the any class ($l=3$). Therefore, the height of this label hierarchy tree is $h=4$.
@@
The ground truth $y_{k}$ for HMC training is derived from this hierarchy. We define it such that more than one label can be correct, but never more than one per hierarchy level $l$. Therefore, the HMC learning rule incorporates multiple correct labels that are learned jointly. The entirety of all classes is denoted as $\lambda$, and $\lambda_{l}$ represents the set of classes at hierarchy level $l$. To prevent the model from exclusively predicting the any class, we assign weights to the superclasses $\mathcal{S}(y)$ of each labeled class $y_{s}$ based on its hierarchy level $l$. Specifically, labels in the leaf nodes $\lambda_{0}$ are assigned the highest weights, while superclasses $\lambda_{1}$ receive lower weights, and so forth. We propose the following formula to weight the new ground truth:
$y_{s}=\begin{cases}\frac{1+l(s)}{h},&\text{if }s\in s\cup\mathcal{S}(y)\\ 0,&\text{otherwise}\end{cases}$ (1)
The altered ground truth $y_{s}$ for class $s$ equals its hierarchy level $l$ if it serves as a superclass $\mathcal{A}$ of the original ground truth label $y$; otherwise, it is assigned 0. According to this ground truth definition, multiple labels assigned to a single measurement trace a path through the hierarchy, reflecting their structural similarity. This implicit relationship modeling enables the model to leverage these relationships for enhanced leaf-node classification as well as meaningful superclass classifications.
@@ III-B Hierarchical Multi-Label Classification
$\mathcal{L}_{HCE}=-\sum_{c\in C}e^{y_{s}}\log\hat{y}_{c}$ (2)
with $m$ being the number of classes in $\lambda$, and $y_{s}$ and $\hat{y}$ being the hierarchically altered ground truth and the model classifications, respectively.

### v3

@@ III-A Data and Label Hierarchy
The label hierarchy serves to group semantically similar classes together. Our proposed label hierarchy for the SemanticKITTI dataset is illustrated in Fig. 2. Each original label is associated with multiple superclasses $\mathcal{S}$ on a level $l$. The dataset’s original ground truth labels $y_{s}$ represent the leaf nodes ($l=0$). Superclasses are structured as follows: the second level aligns with SemanticKITTI’s suggestion ($l=1$), and the third level introduces binary classes, namely static and dynamic ($l=2$). The root node encompasses only the any class ($l=3$). Therefore, the height of this label hierarchy tree is $h=4$.
@@
The ground truth $\eta_{c}$ for HMC training is derived from this hierarchy. We define it such that more than one label can be correct, but never more than one per hierarchy level $l$. Therefore, the HMC learning rule incorporates multiple correct labels that are learned jointly. The entirety of all classes is denoted as $\lambda$, and $\lambda_{l}$ represents the set of classes at hierarchy level $l$. To prevent the model from exclusively predicting the any class, we assign weights to the superclasses $\mathcal{S}(y_{s})$ of each labeled class $y_{s}$ based on its hierarchy level $l$. Specifically, labels in the leaf nodes $\lambda_{0}$ are assigned the highest weights, while superclasses $\lambda_{1}$ receive lower weights, and so forth. We propose the following formula to weight the new ground truth:
$\eta_{c}=\begin{cases}\frac{1+l(c)}{h},&\text{if }c\in y_{s}\cup\mathcal{S}(y_{s})\\ 0,&\text{otherwise}\end{cases}$ (1)
The altered ground truth $\eta_{c}$ is computed from the original subclass $y_{s}$ and its superclasses $\mathcal{S}(y_{s})$. For all other classes it is assigned 0. According to this ground truth definition, multiple labels assigned to a single measurement trace a path through the hierarchy, reflecting their structural similarity. This implicit relationship modeling enables the model to leverage these relationships for enhanced leaf-node classification as well as meaningful superclass classifications.
@@ III-B Hierarchical Multi-Label Classification
$\mathcal{L}_{HCE}=-\sum_{c\in\lambda}e^{\eta_{c}}\log\hat{y}_{c}$ (2)
with $c$ being a given class in $\lambda$, and $\eta_{c}$ and $\hat{y}_{c}$ being the hierarchically altered ground truth and the model classifications for that class, respectively.